Can someone please show all your work and steps fo

Can someone please show all your work and steps for this problem? Please write really legibly and make it easier for me to understand... The answer is ALREADY there so PLEASE make sure to double check your work and make sure your answer matches. Thanks in advance!
Answer #1

There are two components of aceleration, linear and radial
distance covered in one rev = 2*pi*r
and velocity after one rev = v
2*2.01*2*pi*r = v^2 .. (1)
Now adial acc = v^2/r = 2*2.01*2*pi =25.2456
Net acc = sqroot(radial^2 + tangential^2) = 25.325 m/s/s